# De Finetti’s problem with fixed transaction costs and regime switching

### v1

@@ Appendix A Proofs
Therefore, the two functions $\max(\min)_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ is increasing in $\beta$ on $(r^{-}_{j},s^{+}_{j})$.
If there is some $i,j,k\geq 1$ such that $s_{k}^{-}<s_{k}^{+}=t_{j}^{-}\leq t_{j}^{+}=r_{i}^{-}<r_{i}^{+}$, one can derive the desired results by using similar arguments. In this vein, one obtains the increasing property of $\max(\min)_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ in $\beta$ over the interval $[\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),{\omega_{1}}(x_{2})]$ under Case (a). The proof is complete.
@@
$\lim_{\mu_{-}\rightarrow\infty}a_{1}$ $=\lim_{\mu_{-}\rightarrow\infty}\frac{\ln{\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}+\mu_{-}}{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}-\mu_{-}}}}{\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}}{\sigma_{-}^{2}}}=\lim_{\mu_{-}\rightarrow\infty}\frac{2\ln\left(\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}+\mu_{-}\right)-\ln\left(2q\sigma_{-}^{2}\right)}{\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}}{\sigma_{-}^{2}}}=0,$
$\lim_{\mu_{-}\rightarrow\infty}a_{2}$ $=\lim_{\mu_{-}\rightarrow\infty}\frac{\ln\left(\theta_{2}^{+}+\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}+\mu_{-}}{\sigma_{-}^{2}}\right)-\ln\left(\theta_{2}^{+}-\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}-\mu_{-}}{\sigma_{-}^{2}}\right)}{\frac{2\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}}{\sigma_{-}^{2}}}=0,$
@@
[…]
@@
$\liminf_{\mu_{-}\rightarrow\infty}\frac{\zeta(z_{1},z_{2})}{\zeta(\tilde{z}_{1},\tilde{z}_{2})}=\infty.$

### v2

@@ Appendix A Proofs
[…]
If there is some $i,j,k\geq 1$ such that $s_{k}^{-}<s_{k}^{+}=t_{j}^{-}\leq t_{j}^{+}=r_{i}^{-}<r_{i}^{+}$, one can derive the desired results by using similar arguments. In this vein, one obtains the increasing property of $\max(\min)_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ in $\beta$ over the interval $\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),\omega_{1}(x_{2})]$ under Case (a). The proof is complete.
@@
$\lim_{\mu_{-}\rightarrow\infty}a_{1}$ $=\lim_{\mu_{-}\rightarrow\infty}\frac{\ln{\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}+\mu_{-}}{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}-\mu_{-}}}}{\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}}{\sigma_{-}^{2}}}=\lim_{\mu_{-}\rightarrow\infty}\frac{2\ln\left(\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}+\mu_{-}\right)-\ln\left(2q\sigma_{-}^{2}\right)}{\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}}{\sigma_{-}^{2}}}=0,$ (A.39)
$\lim_{\mu_{-}\rightarrow\infty}a_{2}$ $=\lim_{\mu_{-}\rightarrow\infty}\frac{\ln\left(\theta_{2}^{+}+\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}+\mu_{-}}{\sigma_{-}^{2}}\right)-\ln\left(\theta_{2}^{+}-\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}-\mu_{-}}{\sigma_{-}^{2}}\right)}{\frac{2\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}}{\sigma_{-}^{2}}}=0,$ (A.40)
@@
$=$ $\liminf_{\mu_{-}\rightarrow\infty}\frac{\left[(1-c_{+})g^{-}_{q}(0)\theta_{2}^{+}e^{\theta_{2}^{+}{(\tilde{z}_{2}-a)}}+(g^{+}_{q}(0)-c_{+}g^{-}_{q}(0))\theta_{1}^{+}e^{-\theta_{1}^{+}{(\tilde{z}_{2}-a)}}\right](z_{2}-z_{1}-\beta)}{(g^{-}_{q}(0)-c_{-}g^{+}_{q}(0))(e^{\theta^{-}_{2}(z_{2}-a)}-e^{\theta^{-}_{2}(z_{1}-a)})-(1-c_{-})g^{+}_{q}(0)(e^{-\theta^{-}_{1}(z_{2}-a)-e^{\theta^{-}_{1}(z_{1}-a)})}}$
$=$ $\liminf_{\mu_{-}\rightarrow\infty}\left\{\frac{(c_{-}e^{-\theta_{2}^{-}a}+(1-c_{-})e^{\theta_{1}^{-}a})\left[(1-c_{+})\theta_{2}^{+}e^{\theta_{2}^{+}{(\tilde{z}_{2}-a)}}-c_{+}\theta_{1}^{+}e^{-\theta_{1}^{+}(\tilde{z}_{2}-a)}\right](z_{2}-z_{1}-\beta)}{(1-c_{-})e^{(\theta_{1}^{-}-\theta_{2}^{-})a}(e^{\theta_{2}^{-}z_{2}}-e^{\theta_{2}^{-}z_{1}}-e^{-\theta_{1}^{-}z_{2}}+e^{-\theta_{1}^{-}z_{1}})}\right$.
@@
$\liminf_{\mu_{-}\rightarrow\infty}\frac{\zeta(z_{1},z_{2})}{\zeta(\tilde{z}_{1},\tilde{z}_{2})}=\infty.$ (A.46)
@@
Acknowledgments
We sincerely appreciate the anonymous reviewers for their insightful and constructive feedback, which has significantly contributed to the improvement of this manuscript. Wenyuan Wang acknowledges the financial support from the National Natural Science Foundation of China (Nos.: 12571508; 12171405; 11661074).